Write the equation of the line, in a slope-intercept form, that passes through the points (5, -10) and has a slope of -1/5
Answer by subudear(62)

equation of line in slope-intercept form is given by
y = mx + c
where m is slope and c is intercept.
Any line passing through (5, -10) with slope of -1/5 is given by
y - (-10) = (-1/5)(x - 5)
y + 10 = (-1/5)x + 1
y = (-1/5)x - 9